Dayton, Inc. manufactured 15,000 units of product last month and identified the following costs associated with the manufacturing activity.

Variable costs:
Direct materials used $ 1,070,000
Direct labor 2,096,000
Indirect materials and supplies 232,000
Power to run plant equipment 215,000
Fixed costs:
Supervisory salaries 923,000
Plant utilities (other than power to run plant equipment) 292,000
Depreciation on plant and equipment (straight-line, time basis) 142,000
Property taxes on building 198,000


Required:

Unit variable costs and total fixed costs are expected to remain unchanged next month. Calculate the unit cost and the total cost if 21,000 units are produced next month. (Round "Unit costs" to 2 decimal places.)
Answer #1

Answer

  • Total Cost for 21000 units = $ 6,613,200
  • Unit Cost for 21000 units = $ 6613200 / 21000 = $ 314.91 per unit
  • Working

Total Cost for 21000 units

Direct material used

$1,498,000

=1070000*21000/15000

Direct Labor

$2,934,400

=2096000*21000/15000

Indirect material & Supplies

$324,800

=232000*21000/15000

Power to run plant equipment

$301,000

=215000*21000/15000

Total Variable cost

$5,058,200

Supervisory Salaries

$923,000

Plant utilities

$292,000

Depreciation on Plant & equipment

$142,000

Property taxes on building

$198,000

Total Fixed Cost

$1,555,000

Total Cost

$6,613,200